marshae:

The following three proportions can all be used to solve the problem about Caroline's salary. Explain why the solutions to each proportion are the same. 36 4 = x 7 4 36 = 7 x 36 x = 4 7 in words

MiraAngel:

@Shadow

Shadow:

Can you write this out cleanly

stormyfurr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow Can you write this out cleanly \(\color{#0cbb34}{\text{End of Quote}}\) I agree. I can not understand that well

marshae:

36/4=x/7 4/36=4/7 36/x=4/7

marshae:

can u understand

stormyfurr:

I can now

marshae:

ok

MiraAngel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @marshae 36/4=x/7 4/36=4/7 36/x=4/7 \(\color{#0cbb34}{\text{End of Quote}}\) Is it supposed to be 36/4=x/7 4/36=7/x and 36/x=4/7 ??????

Shadow:

Yeah the 2nd one is missing x and has two fours

marshae:

the second one 4/36=7/x

Shadow:

\[\frac{ 36 }{ 4 } = \frac{ x }{ 7 }\] \[\frac{ 4 }{ 36 } = \frac{ 7 }{ x }\] \[\frac{ 36 }{ x } = \frac{ 4 }{ 7 }\]

marshae:

yes

Shadow:

I mean just based off of looking at it, you can see that the same proportions are being represented x/7, 7/x, and in the last one x is across from 7, so similar statements of proportionality are being made. So if you were to solve for x in each of these situations, you would get x = the same # because they're all saying the same thing but in a different way.

marshae:

it's correct

marshae:

thank u guys

MiraAngel:

Great job Shadow. :) I hope he helped! Good luck! If you have any other questions, feel free to reach out to any of us on qc :)